OpenStudy (whiterice):

Find the coordinates of the other endpoint of the​ segment, given its midpoint and one endpoint.​ (Hint: Let​ (x,y) be the unknown endpoint. Apply the midpoint​ formula, and solve the two equations for x and​ y.) midpoint ​(0​,7​), endpoint ​(−2​,0​)

OpenStudy (visarika23):

(x+(-2))/2=0 (y+0)/2=7 SOLVE and get the values of x and y